OpenStudy (anonymous):

Algebra 2 help, please :D

OpenStudy (studygurl14):

what do you need help with?

OpenStudy (anonymous):

The function c(x) = 3x − 4 determines how many cupcakes need to be purchased for a birthday party, where x is the number of kids at the party. The party planner uses d(c(x)) to find the amount of money to bring for the cupcake purchase. The function d(x) = 2x + 5. Solve for how much money to bring when there are 7 kids at the party. 17 19 25 39

OpenStudy (anonymous):

39

OpenStudy (studygurl14):

Okay, so first solve function c(x), with 7 as x. Then solve d(x), but putting the answer of c(x) as x in that equation

OpenStudy (studygurl14):

So... c(x) = 3x - 4 c(7) = 3(7) - 4 c(7) = 21 - 4 c(7) = 17 d(x) = 2x + 5 d(17) = 2(17) + 5 d(17) = 34 + 5 d(17) = 39 39 is the answer :)